OpenStudy (vortish):

factor the trinomial 3x^2+11x-4

OpenStudy (anonymous):

gimme a medal first.

OpenStudy (vortish):

why should i give you a medal first

OpenStudy (anonymous):

becuz u ant an answer

OpenStudy (anonymous):

*want

OpenStudy (vortish):

i would like help not just a answer and i do not like being strong armed

OpenStudy (anonymous):

Factor your on trinomial. RALLOED! You just got Ralloed!

OpenStudy (vortish):

i would like help with this problem just not the answer

OpenStudy (vortish):

learn to spell

OpenStudy (anonymous):

Bye. Nice working with you laddy.

OpenStudy (whpalmer4):

\[3x^2+11x -4\] To factor this, we're going to have \[(3x+a)(x+b) = 3x^2 +ax +3bx+ab = 3x^2 + (a+3b)x + ab\]But we want that to be equal to our original equation, so \[3x^2 + (a+3b)x + ab = 3x^2 + 11x -4\]and therefore\[(a+3b=11)\]\[ab=-4\] Solve those two equations for the values of a and b and substitute them into \[(3x+a)(x+b)\]for the factored polynomial.

OpenStudy (anonymous):

it is (3x-1) ( x+4)